# Visibly Fair Mechanisms

### v1

@@ A.8 More on efficiency
Statement 1: Fix a message $m\in M$. Suppose $a\in\mathcal{A}$ is a visibly fair allocation that is not visibly efficient under $m$. Then there exists another allocation $a^{\prime}\in\mathcal{A}$ such that for all $i\in I$ with $a^{\prime}_{i}\not=a_{i}$, we have $a^{\prime}_{i}\succ_{m_{i}}a_{i}$.
Let $i^{*}$ be the highest ranking officer among the ones that are allocated to a different state under $a$ and $a^{\prime}$ and let $s^{*}$ be her assigned state. That is, for
@@
Allocation $a$ must be visibly unfair for officer $i^{*}$ under $m$. This is because either $a_{i}=s^{*}$ for some $i\in\bar{I}$, or |$|\{i\in I:a_{i}=s^{*}\}|<q_{s^{*}}$. Yet $s^{*}\succ_{m_{i^{*}}}a_{i^{*}}$.
Statement 2: Let $I=\{1,2\}$, $S=\{s_{1},s_{2}\}$, $q_{s_{1}}=q_{s_{2}}=1$, and $\pi(i)=i$. Additionally, assume zonal message space, $Z=\{z_{1}\}$ with $z_{1}=\{s_{1},s_{2}\}$. For preferences ${s_{1}}\succ_{1}s_{2}$ and $s_{1}\succ_{2}{s_{2}}$, the allocation $(a_{1},a_{2})=(s_{2},s_{1})$ is visibly efficient but not visibly fair under the truthful message.
@@
Statement 2 and 3: Let $I=\{1,2\}$, $S=\{s_{1},s_{2},s_{3}\}$, $q_{s_{1}}=q_{s_{2}}=q_{s_{3}}=1$, and $\pi(i)=i$. Additionally, assume a zonal message space, $Z=\{z_{1},z_{2}\}$ with $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. For preferences $s_{3}\succ_{1}{s_{1}}\succ_{1}s_{2}$ and $s_{1}\succ_{2}{s_{3}}\succ_{2}s_{2}$, the allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is visibly fair and visibly efficient under the truthful message as $s_{1}$ and $s_{3}$ are incomparable. However, it is not Pareto efficient. ∎
@@
    A visibly fair allocation under ${m}$ may not be visibly fair under $\hat{m}$.
@@
    A visibly efficient allocation under ${m}$ may not be visibly efficient under $\hat{m}$.
@@
Statement 1: If allocation $a$ is not visibly fair under $m$ because there exist some $i\in I$ such that there is a $j\in I$ such that $a_{i}\not=a_{j}$, $\pi(i)<\pi(j)$, and $a_{j}\succ_{m_{i}}a_{i}$. Since $a_{j}\succ_{\hat{m}_{i}}a_{i}$, $a$ cannot be visibly fair under $\hat{m}$.
@@
If allocation $a$ is not visibly fair under $m$ because there exist some $i\in I$ such that there is a $s\in S$ such that $a_{i}\not=s$, |$|\{i\in I:a_{i}=s\}|\leq q_{s}$, and $s\succ_{m_{i}}a_{i}$. Since $s\succ_{\hat{m}_{i}}a_{i}$, $a$ cannot be visibly fair under $\hat{m}$.
@@
Statement 3: If allocation $a$ is not visibly efficient under $m$, then there exists another allocation $a^{\prime}$ such that for all $i\in I$ with $a^{\prime}_{i}\not=a_{i}$, we have $a^{\prime}_{i}\succ_{m_{i}}a_{i}$. Since $a^{\prime}_{i}\succ_{\hat{m_{i}}}a_{i}$ for all such $i\in I$, $a$ cannot be visibly efficient under $\hat{m}$.
@@
Statement 2 and 4: Let $I=\{1,2\}$, $S=\{s_{1},s_{2},s_{3}\}$, $q_{s_{1}}=q_{s_{2}}=q_{s_{3}}=1$, and $\pi(i)=i$. Additionally, assume a zonal message space, $Z=\{z_{1},z_{2}\}$ with $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. For preferences $s_{3}\succ_{1}{s_{1}}\succ_{1}s_{2}$ and $s_{1}\succ_{2}{s_{3}}\succ_{2}s_{2}$, the allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is visibly fair and visibly efficient under the truthful message $m$ as $s_{1}$ and $s_{3}$ are incomparable.
@@
Consider the zonal message space $Z=\{S\}$. The truthful message $\hat{m}=(\succ_{i})_{i\in I}$ contains more preference information than $m$. However, allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is not visibly fair and visibly efficient under $\hat{m}$. ∎
Appendix B Appendix: Indirect Message Spaces
B.1 Rank-order lists in College Admissions
@@
23 out of 31 provinces in China implement the structured rank-order list system, in which majors are effectively nested under colleges, as noted by Hu et al. (2025). These provinces include: Shanghai, Beijing, Tianjin, Hainan, Jiangsu, Fujian, Hubei, Hunan, Guangdong, Heilongjiang, Gansu, Jilin, Anhui, Jiangxi, Guangxi, Shanxi, Henan, Shaanxi, Ningxia, Sichuan, Yunnan, Tibet, and Xinjiang. For illustrative purposes, we include screenshots of the official college-major preference form from Shanghai.
B.2 Reserve Officer Training Corps (ROTC) Mechanism

### v2

@@ A.8 More on efficiency
This section explores the relationship between visible efficiency and fairness in allocation mechanisms. An allocation is visibly efficient under message profile $m$ if no alternative allocation visibly Pareto dominates it (meaning no reallocation would make all affected agents better off according to their reported preferences). The key results show that visibly fair allocations are always visibly efficient, but the converse doesn’t hold. When comparing visible efficiency to true Pareto efficiency (based on actual preferences), every Pareto efficient allocation is visibly efficient under truthful messages, but visible efficiency doesn’t guarantee Pareto efficiency—even visibly fair allocations can be Pareto inefficient. Furthermore, when message $\hat{m}$ contains more preference information than $m$, allocations that are visibly fair or efficient under $\hat{m}$ remain so under $m$, but not vice versa, demonstrating that visible fairness and efficiency depend critically on the message spaces.
@@
Statement 1: Fix a message profile $m\in M$. Suppose $a\in\mathcal{A}$ is a visibly fair allocation that is not visibly efficient under $m$. Then there exists another allocation $a^{\prime}\in\mathcal{A}$ such that for all $i\in I$ with $a^{\prime}_{i}\not=a_{i}$, we have $a^{\prime}_{i}\succ_{m_{i}}a_{i}$.
Let $i^{*}$ be the highest ranking officer among the ones that are allocated to a different state under $a$ and $a^{\prime}$ and let $s^{*}$ be her assigned state under $a^{\prime}$. That is, for
@@
The allocation $a$ must be visibly unfair under $m$. This is because either $a_{i}=s^{*}$ for some $i\in\bar{I}$, or $|\{i\in I:a_{i}=s^{*}\}|<q_{s^{*}}$. Yet $s^{*}\succ_{m_{i^{*}}}a_{i^{*}}$.
Statement 2: Let $I=\{i_{1},i_{2}\}$, $S=\{s_{1},s_{2}\}$, and $q_{s_{1}}=q_{s_{2}}=1$. Additionally, assume zonal message space for both officers, $Z=\{z_{1}\}$ with $z_{1}=\{s_{1},s_{2}\}$. For preferences ${s_{1}}\succ_{i_{1}}s_{2}$ and $s_{1}\succ_{i_{2}}{s_{2}}$, the allocation $(a_{1},a_{2})=(s_{2},s_{1})$ is visibly efficient but not visibly fair under the truthful message.
@@
Statement 2 and 3: Let $I=\{i_{1},i_{2}\}$, $S=\{s_{1},s_{2},s_{3}\}$, and $q_{s_{1}}=q_{s_{2}}=q_{s_{3}}=1$. Additionally, assume a zonal message space, $Z=\{z_{1},z_{2}\}$ with $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. For preferences $s_{3}\succ_{i_{1}}{s_{1}}\succ_{i_{1}}s_{2}$ and $s_{1}\succ_{i_{2}}{s_{3}}\succ_{i_{2}}s_{2}$, the allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is visibly fair and visibly efficient under the truthful message, as $s_{1}$ and $s_{3}$ are not comparable. However, it is not Pareto efficient. ∎
@@
    A visibly fair allocation under m{m} may not be visibly fair under $\hat{m}$.
@@
    A visibly efficient allocation under m{m} may not be visibly efficient under $\hat{m}$.
@@
Statement 1: There are two cases: (i) Allocation $a$ is not visibly fair under $m$ because there exist some $i\in I$ such that there is a $j\in I$ such that $a_{i}\not=a_{j}$, $\pi(i)<\pi(j)$, and $a_{j}\succ_{m_{i}}a_{i}$. Since $_{j}\succ_{\hat{m}_{i}}a_{i}, aa$ cannot be visibly fair under $\hat{m}$.
@@
(ii) Allocation $a$ is not visibly fair under $m$ because there exist some $i\in I$ such that there is a $s\in S$ such that $a_{i}\not=s$, $|\{i\in I:a_{i}=s\}|\leq q_{s}$, and $s\succ_{m_{i}}a_{i}$. Since $s\succ_{\hat{m}_{i}}a_{i}$, $a$ cannot be visibly fair under $\hat{m}$.
@@
Statement 3: We use the contrapositive. If allocation $a$ is not visibly efficient under $m$, then there exists another allocation $a^{\prime}$ such that for all $i\in I$ with $a^{\prime}_{i}\not=a_{i}$, we have $a^{\prime}_{i}\succ_{m_{i}}a_{i}$. Since $a^{\prime}_{i}\succ_{\hat{m_{i}}}a_{i}$ for all such $i\in I$, $a$ cannot be visibly efficient under $\hat{m}$.
@@
Statement 2 and 4: Let $I=\{i_{1},i_{2}\}$, $S=\{s_{1},s_{2},s_{3}\}$, and $q_{s_{1}}=q_{s_{2}}=q_{s_{3}}=1$. Additionally, assume a zonal message space for every officer, $Z=\{z_{1},z_{2}\}$ with $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. For preferences $s_{3}\succ_{i_{1}}{s_{1}}\succ_{i_{1}}s_{2}$ and $s_{1}\succ_{i_{2}}{s_{3}}\succ_{i_{2}}s_{2}$, the allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is visibly fair and visibly efficient under the truthful message $m$ as $s_{1}$ and $s_{3}$ are not comparable.
@@
Consider the zonal message space $Z=\{S\}$. The truthful message $\hat{m}=(\succ_{i})_{i\in I}$ contains more preference information than $m$. However, allocation $(a_{1},a_{2})=(s_{1},s_{3})$ is not visibly fair nor visibly efficient under $\hat{m}$. ∎
Appendix B Appendix: Proofs that examples satisfy sequential solvency
Proof that Example 5.2 satisfies sequential solvency.
Assume, for contradiction, that there exists an officer $i$ who, given some feasible placement $a_{-i}$ of the other $7$ officers, has no admissible state. Without loss of generality let $t_{i}=R_{1}$.
  • If any seat in $R_{2}$ were vacant it would be admissible for $i$ (because the bound $(R_{1},2)$ does not apply there), contradicting the assumption. Hence both states in $R_{2}$ are fully occupied ($4$ officers).
  • The local cap in $R_{2}$ is $(R_{2},2)$; therefore at most two of the four occupants can be of type $R_{2}$. So at least 42=24-2=2 of them are of type $R_{1}$.
  • Among the remaining $7$ officers, at most one $R_{1}$-type can still be in $R_{1}$ (since two are already in $R_{2}$). Thus $(R_{1},2)$ is not binding in $R_{1}$.
  • Only $7$ officers are placed, but $R_{1}$ has capacity $q_{s_{1}}+q_{s_{2}}=4$; hence at least one seat in $R_{1}$ is empty.
But then, $(R_{1},2)$ is not binding in $R_{1}$ and there is spare capacity in $R_{1}$, contradicting $i$ not having an admissible state.
Proof that Example 5.3 satisfies sequential solvency.
Suppose, for contradiction, that some type–1 doctor $i$ and feasible placement $a_{-i}$ of the other $26$ doctors leave $i$ without an admissible hospital. This can only happen if:
  1. (i)
    all hospitals in $R_{2}$ and $R_{3}$ are full; or
  2. (ii)
    the only vacancies there are urban and the urban cap ($19$) is binding.
Case (i): $Q(R_{2})+Q(R_{3})=24$, so at most $24$ of the others are outside $R_{1}$. Then at most $2$ are in $R_{1}$, leaving $\geq 10$ free seats there; the local cap $(R_{1},6)$ is not binding, so $i$ could be placed in $R_{1}$.
Case (ii): With $19$ urban doctors, the remaining 2619=726-19=7 occupy rural seats, leaving 127=512-7=5 rural vacancies overall. If all rural seats in $R_{2}$ and $R_{3}$ were full, they would require $8$ rural occupants, contradicting the total of $7$. Thus some rural hospital in $R_{2}$ or $R_{3}$ has a vacancy admissible to $i$.
In both cases we contradict the assumption; hence sequential solvency holds.
Appendix C Appendix: Indirect Message Spaces
C.1 Preference Elicitation in All India Services
The 2017 Cadre Allocation Policy for India’s All India Services, including the Indian Administrative Service (IAS), Indian Police Service (IPS), and Indian Forest Service (IFoS), introduces a zonal system that divides all states and union territories into five geographical zones, requiring candidates to first indicate their zone preferences in descending order, followed by cadre preferences within each preferred zone. For illustrative purposes, we include a screenshot of the submitted preferences from 2020 IFoS examination.
C.2 Rank-order lists in Chinese College Admissions
@@
23 out of 31 provinces in China implement the structured rank-order list system, in which majors are effectively nested under colleges, as noted by Hu et al. (2025). These provinces include: Shanghai, Beijing, Tianjin, Hainan, Jiangsu, Fujian, Hubei, Hunan, Guangdong, Heilongjiang, Gansu, Jilin, Anhui, Jiangxi, Guangxi, Shanxi, Henan, Shaanxi, Ningxia, Sichuan, Yunnan, Tibet, and Xinjiang. For illustrative purposes, we include a screenshot of the official college-major preference form from Shanghai.
C.3 Reserve Officer Training Corps (ROTC) Mechanism